OpenStudy (lopus):

purchased monthly 1000 g of food for birds. bird "A" eats 200 g more than bird "B". How much each bird eats?

OpenStudy (anonymous):

b + 200 g = A A + B = 1000g Can you see it better now?

OpenStudy (lopus):

a-b=200 a+b=1000 ---------------- 2a=1200 a=600 b=400 i'm right?

OpenStudy (anonymous):

I am not following your process, can you show it?

OpenStudy (lopus):

only sum equations

OpenStudy (anonymous):

what?

OpenStudy (lopus):

just add the equations

OpenStudy (lopus):

i get it, thanks